# Diffusion MRI with Machine Learning

### v1

@@ III-A Validation
  • The methods are usually trained and validated on normal brains. This is a major flaw because brain pathologies may represent vastly different and heterogeneous macro-structure and drastically alter the dMRI signal [127, 337, 338]. Brain pathologies such as tumors can also drastically alter the brain anatomy and impact the performance of automatic tract analysis methods [195, 194]. For micro-structure mapping, some studies have presented preliminary validation on selected pathologies such as MS [26] and stroke [38]. One study trained a model on healthy brain data to jointly estimate the DTI, DKI, and multi-compartment models and tested the model on stroke patients [71]. Results showed that the model accuracy on pathological tissue suffered more for multi-compartment models (e.g., NODDI) than for DTI. Similarly, a few studies have validated tract-specific analysis methods on clinical data. For example, one study [339] showed that automatic tract segmentation with TractSeg [27] was more accurate in detecting micro-structural changes due to traumatic brain injury than a method that computed DTI followed by manual tract segmentation. Lie at al. [197] validated their joint tract segmentation and registration on a clinical population. Although most of the published results on abnormal brains are positive and encouraging, they are not sufficient. Validation of deep learning methods on abnormal brain data is especially important because these methods are highly complex and difficult to interpret, and because the generalizability of these methods to unseen data is hard to predict [340]. As an example, for micro-structure mapping one study has found that training with a loss function in terms of error in micro-structure estimation may lead to better generalizability to pathological brains than loss formulations in terms of signal reconstruction error [85]. Other factors may also influence the generalizability of machine learning / deep learning methods to pathological test data, but there has been very little work to investigate this topic.
@@
  • Validation on external datasets is also necessary. A number of works have reported successful application of trained machine learning models on data from other scanners/centers without any data harmonization or re-training [73]. Other studies have suggested that when dMRI intensities or subject demographics deviate from training data, intensity normalization, data harmonization, or model re-training/fine-tuning may be necessary [26, 341]. However, data harmonization in dMRI is a complex problem of its own as we have briefly touched on in Section II-H. Moreover, we think there may be a tendency in the published works not to report discouraging results when it comes to validation on external data. Given the significant diversity across datasets and applications and strong dependence of machine learning methods on training data distribution, model settings, and design choices, more extensive validations are needed to assess the true advantages of machine learning methods in real-world settings with heterogeneous data.
@@
III-B Inherent limitations and drawbacks of machine learning approaches
Despite the indisputable advantages of machine learning methods, some of which have been discussed in Section I-C, machine learning methods also suffer from inherent limitations. We describe some of these below, with a focus on applications in dMRI analysis.
@@ III-B Inherent limitations and drawbacks of machine learning approaches
  • Generalizability of machine learning methods to new test data can be affected by factors that may not be clear beforehand [67]. For instance, it has been suggested that factors such as signal intensity and acquisition parameters such as echo time may impact the performance of a neural network model to the extent that a complete re-training may be necessary [344, 26]. Performance of conventional (i.e., non-learning) methods is probably much less affected by such unforeseen factors.
@@
  • Closely related to the above points is the issue of out-of-distribution (OOD) data, which refers to data points that are outside the distribution of the training data. Detecting OOD data in higher-dimensional signal spaces is very challenging, and machine learning methods have no performance guarantee on OOD data [345, 346]. A machine learning technique that works well on in-distribution data can experience unexpected and catastrophic failure on OOD test data. Obtaining training data that is sufficiently rich and heterogeneous to represent all data that will be potentially encountered at test time can be difficult if not impossible because many factors may influence the data and label distributions. In tractography, for example, data noise and artifacts can be important factors and the trained model may produce inferior results if the training and test data are different with regard to these aspects [22, 151]. Non-learning methods are less prone to such failures.
@@
  • Most modern machine learning methods are complex and difficult to interpret, a topic that is further discussed below in Section III-E. This means that they may suffer from shortcomings that are not easy to recognize. For example, it has been shown that machine learning methods can lead to systematic and unpredictable biases in micro-structure mapping from dMRI signal [53, 67]. It has also been shown that this bias can depend on factors such as SNR, partial volume effects, and distribution of the micro-structure values in the training and test data [53, 67]. As a result, predictions of these methods may lead to erroneous conclusions when studying the impact of pathology on micro-structure [67]. While such biases may be present in many machine learning methods, discovering them requires extensive validation and analysis that is missing in most published works.
III-C Modeling considerations
III-C1 Data representation
The dMRI data used as input to machine learning models are typically in the so-called q space, where each of the measurements in a voxel is associated with a gradient strength and a gradient direction. The gradient strength/direction are typically different between subjects/scans. Two approaches that are commonly used to create a uniform representation of the data include: (1) interpolation unto a fixed spherical grid, where a typical grid size is approximately 100 [39], and (2) representation in spherical harmonics. There exists several formulations for extending the spherical harmonics representation to multi-shell data [347, 348]. Zucchelli et al. [349, 45] have advocated for signal representation in terms of their rotationally invariant features [350]. They have shown that, compared with the two common representations mentioned above, their representation leads to better estimation of multi-compartment models [349] and FOD [350]. Another notable example is sparse representation in dictionaries [61]. In [64], dictionaries that were separable in spatial and angular domains [351] were used. Compared with non-separable dictionaries, separable formulations significantly reduce the number of parameters and may improve the analysis results [64]. Another flexible representation framework is based on graphs. This framework was successfully used for dMRI signal interpolation in [31], where the authors constructed the adjacency matrix in terms of the difference between pairs of measurements. They computed the difference in terms of spatial distance, angular difference between gradient directions, and difference in gradient strengths. A similar representation was used in [313] to extend the non-local means denoising algorithm to dMRI. Another approach, proposed in [82], consists of projecting the q-space data onto the standard (xy, xz, and yz) planes and binning them into fixed grids. For multi-shell data, each shell is binned separately. For applications that require large spatial fields of view, such as in tract segmentation [27], limited GPU memories typically do not allow for using the dMRI data as model input. Instead, as mentioned above in Sections II-C, II-D, and II-E, DTI or FOD estimations may be computed and used as input for machine learning models.
III-C2 Neural network architecture
@@
Standard convolutional networks are equivariant to translation, which is one of the key reasons for their success in computer vision. However, they are not equivariant to rotation. Rotation-invariance and rotation-equivariance (i.e., a rotation of the input to result in an identical rotation in the output) are highly desired in many applications. Incorporating these properties in the network architecture removes the need for the model to learn every possible orientation of the object/scene, which would otherwise require larger datasets or massive data augmentation. It has been shown to improve the performance in various computer vision tasks [354]. In dMRI, the q-space data in each voxel can be represented as spherical maps. Therefore, in addition to the rotations in the voxel (physical) space, q-space rotations should also be accounted for [110]. Most standard neural network models such as CNNs have been designed for signals in Euclidean space such as digital images. An ideal convolutional operation for q-space data should be either rotation-invariant (e.g., for estimating scalar maps of micro-structure) or rotation-equivariant (e.g., for FOD estimation). Moreover, it should allow for irregular sampling on the sphere.
@@
Muller et al. [355] have argue that standard (i.e., non-rotationally-equivariant) networks, when applied in dMRI, are at the risk of both under-fitting and over-fitting. This is because they need to learn the rotation-equivariance relationships from training data, which may be biased towards certain orientations. Incorporating rotation-equivariance will eliminate the need for learning these constraints, enable better weight sharing, improve generalization to orientations not seen in training data, and allow the model to concentrate on learning the actual function of interest [109, 355].
@@
Various techniques have been proposed to build rotation-equivariance into neural networks and to extend standard convolutions to signals on the sphere. We briefly explain some of these below.
@@
  • Some studies have proposed to discretize the surface of the sphere and apply standard 2D convolutions on the resulting grid [356, 357, 358]. A straightforward discretization can be achieved by using an equiangular spacing in the standard spherical-polar coordinate system. Boomsma and Frellsen [357] have suggested that a better discretization can be achieved using the cubed-sphere representation originally developed by [359]. In this representation, each point on the sphere is represented by the coordinates in one of the six patches of a circumscribed cube. They define cubed-sphere convolution as applying a standard 2D convolution on a uniformly spaced grid on the six sides of the cube. This method can be extended to concentric shells in a ball.
@@
  • A common approach in computer vision applications is to divide the surface of the sphere into small regions and process each region as a planar signal with standard 2D convolutions [360, 361]. This processing is usually followed by hierarchical pooling. However, this approach does not result in rotation-equivariance.
@@
Numerous works have reported that these specialized network architectures can improve the performance of deep learning models in different dMRI analysis tasks [364, 366, 273, 367, 368]. Bouza et al. found that their proposed convolution operation improved the state of the art in classification of Parkinson’s Disease patients versus healthy controls and in FOD reconstruction [364]. Sedlar et al. reported improved estimation of NODDI parameters with a rotationally-invariant CNN [369], while Liu et al. show that their network results in a 10% reduction in the number of trainable weights without a reduction in performance in a tissue segmentation application [370]. Kerkela et al. [367] found that a rotation invariant spherical CNN performed better than a standard MLP in estimating a multi-compartment model. For angular super-resolution of dMRI data, Lyon et al. showed that a parametric continuous convolution network outperformed a standard CNN while reducing the number of parameters by an order of magnitude [274]. Other dMRI study have shown that rotationally-equivariant networks can lead to more accurate tissue segmentation and lesion segmentation compared with standard networks [368, 355]. Goodwin-Allcock et al. showed that including rotation equivariance could improve the generalizability of deep learning models to new gradient schemes and reduce the required training data size [366].
III-C3 Loss functions
@@
For tissue micro-structure mapping, one study compared a training loss function based on the error in micro-structure prediction and another in terms of the error in signal prediction [85]. The former is by far more common in machine learning-based micro-structure mapping works, whereas the latter is the common loss function in standard optimization-based methods. It was shown that although the former led to lower micro-structure prediction errors when the noise was modeled accurately, the latter resulted in lower signal reconstruction error and it was more practical because it eliminated the need for reference micro-structure and enabled training with synthetic and real data [85]. The authors speculated that the weakness of a loss function based on signal reconstruction error could be due to the fact that neural networks might tend to learn features that originate from the noise floor. In the context of IVIM model fitting, a similar method of training based on the signal prediction error was proposed in [34], where the authors used the terms “unsupervised” and “physics-informed” to describe their method. Parker et al. argue that the use of root mean square loss function is not consistent with the measurement noise distribution in dMRI [88]. They propose a loss based on Rician likelihood and show that, for estimation of apparent diffusion coefficient (ADC) and IVIM parameters, it leads to lower estimation bias at lower SNR.
@@
The quest for a proper loss function in micro-structure estimation has also been pursued in [92], where the authors argue that loss functions based purely on the predicted signal are inadequate. Instead, they propose two alternative loss functions that additionally incorporate the error in micro-structure estimation. One of their loss functions consists of the $\ell_{1}$ norms of the errors of the dMRI signal and the micro-structure indices. The other loss function, which they name the spherical variance loss, is based on the spherical variance of the dMRI signal that is correlated with micro-structure [371]. Experiments show that, compared with standard loss functions such as the $\ell_{1}$, $\ell_{2}$, and Huber loss of the signal prediction error, these two augmented loss functions result in more accurate prediction of a range of parameters including GFA, DKI, and NODDI. However, they do not compare these new loss functions with a loss function based on the error in predicted micro-structure index.
III-D Data, ground truth, and evaluation
@@
For many dMRI applications, obtaining data with reliable target/label is difficult or impossible. In some applications, such as those involving segmenting or analyzing specific white matter tracts, labels can be obtained from human experts. However, such labels are known to suffer from significant inter- and intra-rater variability [171, 156]. In certain applications, histological data can be used to establish a reasonable ground truth, but at considerable cost. For estimation of intra-axonal water exchange time, Hill et al. used histology analysis of an in-vivo mouse model of demyelination using electron microscopy to generate ground truth [43]. Specifically, they measured the thickness of the myelin sheath as a measure of exchange time. Surprisingly, they found that the prediction of their random forest regression model based on dMRI signal was strongly correlated with histological measurements, with a Pearson correlation coefficient of 0.98. Nath et al. used confocal histology to generate data for FOD estimation [111]. Since the data obtained with this technique was costly and very limited, they performed massive data augmentation to increase the effective size of their data. For data harmonization, a large purely dMRI scan-rescan dataset was used to boost a small dataset of paired histology and dMRI [298, 301]. Another work used co-registered dMRI and histology data to develop a machine learning method for estimating several micro-structural indices and showed that it achieved higher estimation accuracy than standard methods [21]. For validating a dMRI-based tract segmentation methods, Lucena et al. [195, 194] have advocated for using navigated transcranial magnetic stimulation to establish a ground truth.

### v2

@@ V-A1 Validation on abnormal/pathological data
For tissue microstructure mapping, some studies have validated the new methods on selected pathologies such as MS [16] and stroke [55]. One study trained a model on healthy brain data to jointly estimate the DTI, DKI, and multi-compartment models and tested the model on stroke patients [165]. Results showed that the model accuracy on pathological tissue suffered more for multi-compartment models (e.g., NODDI) than for DTI. Similarly, a few studies have validated tract-specific analysis methods on clinical data. For example, one study [363] showed that automatic tract segmentation with TractSeg [17] was more accurate in detecting microstructural changes due to traumatic brain injury than a method based on manual tract segmentation. Li at al. [308] validated their joint tract segmentation and registration method on a clinical population. Although most of the published results on abnormal brains are positive and encouraging, they are not sufficient. Validation of deep learning methods on abnormal brain data is especially important because these methods are highly complex and difficult to interpret, and because the generalizability of these methods to unseen data is hard to predict [364]. As an example, for tissue microstructure mapping one study has found that training with a loss function in terms of error in microstructure estimation may lead to better generalizability to pathological brains than loss functions in terms of signal reconstruction error [178]. Other factors may also influence the generalizability of machine learning methods to pathological test data, but there has been little work to investigate this topic.
V-A2 Validation on external data
@@
A number of works have reported successful application of trained machine learning models on data from other scanners/centers without any data harmonization or re-training [162]. Other studies have suggested that when dMRI intensities or subject demographics deviate from training data, intensity normalization, data harmonization, or model re-training/fine-tuning may be necessary [16, 365]. However, data harmonization in dMRI is a complex problem of its own as mentioned on in Section IV-C. Moreover, there may be a tendency in the published works not to report discouraging results when it comes to validation on external data. Given the significant diversity across datasets and applications and strong dependence of machine learning methods on training data distribution, model settings, and design choices, more extensive validations are needed to assess the true advantages of machine learning methods in real-world settings with heterogeneous data.
V-A3 Challenges of performing fair and extensive validations and comparisons
@@
V-B Inherent limitations and drawbacks of machine learning approaches
Despite the indisputable advantages of machine learning methods, some of which have been discussed in Section III, machine learning methods also suffer from inherent limitations. This section describes some of these, with a focus on applications in dMRI analysis.
V-B1 Dependence on the training data distribution
@@ V-B2 Unpredictable generalizability to unseen test data
Generalizability of machine learning methods to new test data can be affected by factors that may not be clear beforehand [168]. For instance, it has been suggested that factors such as signal intensity and acquisition parameters such as echo time may impact the performance of a neural network model to the extent that a complete re-training may become necessary [368, 16]. Performance of conventional (i.e., non-learning) methods is probably much less affected by such unforeseen factors.
@@
Closely related to the above point is the issue of out-of-distribution (OOD) data, which refers to data samples that are outside the distribution of the training data. Detecting OOD data in higher-dimensional signal spaces is very challenging, and machine learning methods have no performance guarantee on OOD data [369, 370]. A machine learning technique that works well on in-distribution data can experience unexpected and catastrophic failure on OOD test data. Obtaining training data that is sufficiently rich and heterogeneous to represent all data that will be potentially encountered at test time can be difficult if not impossible because many factors may influence the data distribution. In tractography, for example, data noise and artifacts can be important factors and the trained model may produce inferior results if the training and test data are different with regard to these aspects [18, 237]. Non-learning methods are less prone to such failures.
@@
It has been shown that machine learning methods can lead to systematic and unpredictable biases in microstructure mapping from dMRI signal [147, 168]. Moreover, this bias can depend on factors such as SNR, partial volume effects, and distribution of the microstructure values in the training and test data [147, 168]. As a result, predictions of these methods may lead to erroneous conclusions when studying the impact of pathology on microstructure [168]. While such biases may be present in many machine learning methods, discovering them requires extensive validation and analysis that is missing in most published works. These suggest that machine learning methods may suffer from shortcomings that are not easy to recognize. In general, most modern machine learning methods are complex and difficult to interpret, a topic that is further discussed below in Section V-E.
V-C Modeling considerations
This section describes some of the technical considerations that commonly arise in developing machine learning models in dMRI.
V-C1 Data representation
The dMRI data used as input to machine learning models are typically in the so-called q space, where each measurement in a voxel is associated with a gradient strength and a gradient direction. The gradient strength/direction are typically different between subjects/scans. Many different approaches have been used to uniformly represent the data.
  • A simple approach is to interpolate the data unto a fixed spherical grid, where a typical grid size is approximately 100 [56].
  • A common approach is representation in spherical harmonics. There exists several formulations for extending the spherical harmonics representation to multi-shell data [371, 372].
  • Zucchelli et al. [373, 62] have advocated for signal representation in terms of their rotationally invariant features [374]. They have shown that, compared with the two common representations mentioned above, their representation leads to better estimation of multi-compartment models [373] and FOD [374].
  • Another approach is sparse representation in dictionaries [158]. In [160], dictionaries that were separable in spatial and angular domains [375] were used. Compared with non-separable dictionaries, separable formulations significantly reduce the number of parameters and may improve the analysis results [160].
  • Graphs offer a flexible representation framework that has been successfully used by prior works. In [49], the authors constructed the adjacency matrix in terms of the difference between pairs of measurements. They computed the difference in terms of spatial distance, angular difference between gradient directions, and difference in gradient strengths. A similar representation was used in [70] to extend the non-local means denoising algorithm to dMRI.
  • Another approach, proposed in [149], consists of projecting the q-space data onto the standard (xy, xz, and yz) planes and binning them into fixed grids. For multi-shell data, each shell is binned separately.
  • For applications that require large spatial fields of view, such as in tract segmentation [17], limited GPU memory typically does not allow for using the dMRI data as model input. In such applications, as mentioned above in Sections IV-F, IV-G, and IV-I, DTI or FOD maps may be computed and used as input for machine learning models.
V-C2 Neural network architecture
@@
Standard convolutional operations are equivariant to translation, which is one of the key reasons for their success in computer vision. However, they are not equivariant to rotation. Rotation-invariance and rotation-equivariance (i.e., a rotation of the input to result in an identical rotation in the output) are desired in many applications. Incorporating these properties in the network architecture removes the need for the model to learn every possible orientation, which would otherwise require large datasets or massive data augmentation. It can improve the performance in various computer vision tasks [378]. In dMRI, the q-space data in each voxel can be represented as spherical maps. Therefore, in addition to the rotations in the voxel (physical) space, q-space rotations should also be accounted for [195]. Most standard neural network models such as CNNs have been designed for signals in Euclidean space such as digital images. An ideal convolutional operation for q-space data should be either rotation-invariant (e.g., for estimating scalar maps of tissue microstructure) or rotation-equivariant (e.g., for FOD estimation). Moreover, it should allow for irregular sampling on the sphere.
@@
Muller et al. [379] have argued that standard (i.e., non-rotationally-equivariant) networks, when applied in dMRI, are at the risk of both under-fitting and over-fitting. This is because they need to learn the rotation-equivariance relationships from training data, which may be biased towards certain orientations. Incorporating rotation-equivariance will eliminate the need for learning these constraints, enable better weight sharing, improve generalization to orientations not seen in the training data, and allow the model to concentrate on learning the actual underlying function of interest [194, 379].
@@
Various techniques have been proposed to build rotation-equivariance into neural networks and to extend standard convolutions to signals on the sphere. Some of these designs are described below.
@@
  • Some studies have proposed to discretize the surface of the sphere and to apply standard 2D convolutions on the resulting grid [380, 381, 382]. A straightforward discretization can use an equiangular spacing in the standard spherical-polar coordinate system. Boomsma and Frellsen [381] suggest that a better discretization can be achieved by using the “cubed sphere” representation originally developed by [383]. In this representation, each point on the sphere is represented by the coordinates in one of the six patches of a circumscribed cube. They define cubed-sphere convolution as applying a standard 2D convolution on a uniformly spaced grid on the six sides of the cube. This method can be extended to concentric shells.
@@
  • A common approach in computer vision is to divide the surface of the sphere into small regions and process each region as a planar signal with standard 2D convolutions [384, 385]. This processing is usually followed by hierarchical pooling. However, this approach does not result in rotation-equivariance.
@@
Numerous works have reported that these specialized network architectures can improve the performance of deep learning models in different dMRI analysis tasks [388, 390, 97, 391, 392]. Bouza et al. found that their proposed convolution operation improved the state of the art in classification of Parkinson’s Disease patients versus healthy controls and in FOD reconstruction [388]. Sedlar et al. reported improved estimation of NODDI parameters with a rotationally-invariant CNN [393], while Liu et al. showed that their network resulted in a 10% reduction in the number of trainable weights without a reduction in performance in a tissue segmentation application [394]. Kerkela et al. [391] found that a rotation invariant spherical CNN performed better than a standard MLP in estimating a multi-compartment model. For angular super-resolution of dMRI data, Lyon et al. showed that a parametric continuous convolution network outperformed a standard CNN while reducing the number of parameters by an order of magnitude [98]. Other dMRI studies have shown that rotationally-equivariant networks can lead to more accurate tissue segmentation and lesion segmentation compared with standard networks [392, 379]. Goodwin-Allcock et al. have shown that including rotation equivariance improves the generalizability of deep learning models to new gradient schemes and reduces the required training data size [390].
V-C3 Optimization objective functions
@@
For tissue microstructure mapping, one study compared a training loss function based on the error in microstructure prediction with another in terms of the error in signal prediction [178]. The former has been by far more common in machine learning methods, whereas the latter is the common loss function in standard optimization-based methods. It was shown that although the former led to lower microstructure prediction errors when the noise was modeled accurately, the latter resulted in lower signal reconstruction error and it was more practical because it eliminated the need for reference microstructure and enabled training with synthetic and real data [178]. The authors speculated that the weakness of a loss function based on signal reconstruction error could be due to the fact that neural networks might tend to learn features that originate from the noise floor. In the context of IVIM model fitting, a similar method of training based on the signal prediction error was proposed in [52], where the authors used the terms “unsupervised” and “physics-informed” to describe their method. Parker et al. argue that the use of root mean square loss function is not consistent with the measurement noise distribution in dMRI [181]. They propose a loss based on Rician likelihood and show that, for estimation of apparent diffusion coefficient (ADC) and IVIM parameters, it leads to lower estimation bias at lower SNR.
@@
Another work on tissue microstructure estimation has argued that loss functions based purely on the predicted signal are inadequate [95]. Instead, the authors propose two alternative loss functions that additionally incorporate the error in microstructure estimation. One of their loss functions consists of the $\ell_{1}$ norms of the errors of the dMRI signal and the microstructure indices. The other loss function, which they name the spherical variance loss, is based on the spherical variance of the dMRI signal that is correlated with microstructure [395]. Experiments show that, compared with standard loss functions such as the $\ell_{1}$, $\ell_{2}$, and Huber loss of the signal prediction error, these two augmented loss functions result in more accurate prediction for a range of parameters including GFA, DKI, and NODDI. However, they do not compare these new loss functions with a loss function based on the error in the predicted tissue microstructure index.
V-D Data, ground truth, and evaluation
For many dMRI applications, obtaining data with reliable target/label is difficult or impossible.
V-D1 Synthesizing data from high-quality dMRI scans
Most commonly, high-quality dMRI measurements have been used to generate the data needed for developing and validating machine learning methods. For tract segmentation/registration, Li et al. used an automatic method based on probabilistic tractography and atlas-based segmentation to generate their training labels [23]. For tissue microstructure mapping and for FOD estimation, most studies have used dense q-space data and estimated ground truth parameters using standard methods [162, 203, 204, 209, 202, 209].
V-D2 Histology/microscopy
@@
In certain applications, histological data can be used to establish a reasonable ground truth, but at considerable cost. For estimation of intra-axonal water exchange time, Hill et al. used histology analysis of an in-vivo mouse model of demyelination using electron microscopy to generate ground truth [60]. Specifically, they measured the thickness of the myelin sheath as a measure of exchange time. Surprisingly, they found that the prediction of their random forest regression model based on dMRI signal was strongly correlated with histological measurements, with a Pearson correlation coefficient of 0.98. Nath et al. used confocal histology to generate data for FOD estimation [196]. Since the data obtained with this technique was costly and very limited, they performed massive data augmentation to increase the effective size of their data. For data harmonization, a large purely dMRI scan-rescan dataset was used to boost a small dataset of paired histology and dMRI [123]. Another work used co-registered dMRI and histology data to develop a machine learning method for estimating several microstructural indices and achieved higher estimation accuracy than standard methods [42]. For validating a dMRI-based tract segmentation methods, Lucena et al. [306, 305] have advocated for using navigated transcranial magnetic stimulation to establish a ground truth.
V-D3 Annotations from human experts
In some applications, such as those involving segmenting or analyzing specific white matter tracts, labels can be obtained from human experts. This approach, however, may suffer from difficulty of defining anatomical ground truth [54], high intra- and inter-expert variability [263, 246], and labeling errors. Such label errors need to be properly accounted for using advanced machine learning techniques [357]. In applications such as tractography and tract analysis, obtaining labels from multiple experts may be useful but costly [265, 18, 237].
V-D4 Data simulation